OpenStudy (anonymous):

Use basic trigonometric identities to simplify the expression: 3sin3 (x) csc(x) + cos2 (x) + 2cos (−x) cos (x) = ?

OpenStudy (anonymous):

is it 3sin^3(x) or 3sin(3x)?

OpenStudy (anonymous):

oh i forgot the carets its 3sin^3 (x) csc(x) + cos ^2 (x) +2cos (-x) cos (x)

OpenStudy (anonymous):

lets start with the first term 3sin^3 (x) csc(x)

OpenStudy (anonymous):

csc(x) is the same as 1/sinx

OpenStudy (anonymous):

yess so it would be like 3sin^3 (x) 1/sinx

OpenStudy (anonymous):

ya, so what would that simplify to?

OpenStudy (anonymous):

2sin^3?

OpenStudy (anonymous):

no, it should be 3sin^2x

OpenStudy (anonymous):

one of the sinx will cancel

OpenStudy (anonymous):

ohh okay

OpenStudy (jdoe0001):

\(\bf 3sin^3(x) csc(x) + cos^2 (x) + 2cos (-x) cos (x)\qquad \begin{cases} csc(x)=\frac{1}{sin(x)}\\ cos(-x)=cos(x) \end{cases} \\ \quad \\ 3\cancel{sin^3(x)}\cdot {\color{brown}{ \cfrac{1}{\cancel{sin(x)}} }} + cos^2 (x) + 2cos ({\color{brown}{ x}}) cos (x) \\ \quad \\ 3sin^2(x)+cos^2(x)+2cos(x)cos(x) \\ \quad \\ \textit{what's }cos(x)cos(x)=?\)